X~exp(λ) with λ=1

1) define Y= X^1/2. Find the support of Y and its density.

2) define Z = X^2 + 2X. Find the support of Z and its density.
Answer #1

Given to eap (d)fa=! So f(x)=ě x 220 2 Now Flag a PCX<X) = fitdt et SI-ena F(X) = 1-e? 0 we have to find the density of Y= x²Given z = x2+2x., So G (20 = P(Z <z) ap(*72x sz) EPL x72x+1.32+1) = platu?< 2+1] = P(-2+0 < *+13MZF P[-2+1 -1 5*<V2+ - 1] PE-G(Z) = (W211 -1) Now 962) = d G (2) dz d -dZ+1 -1) [1-6 1 ) tzt1-1) 이 e dz let t=&zt 1-1 then 912) - d dt -t.dt dz e-t dd2+1-